Find the orthogonal canonical reduction of the quadratic form
-x²+y²+z²+xy+xz-6yz
. Also, find its principal axes.
1
Expert's answer
2020-03-19T17:33:16-0400

First, determine the matrix of the quadratic form "Q = -x^2 + y^2 + z^2 + xy + xz - 6yz".


"A = \begin{bmatrix}
 -1 & \frac{1}{2} & \frac{1}{2} \\
 \frac{1}{2} & 1 & -3 \\
 \frac{1}{2} & -3 & 1
\end{bmatrix}"


Find an eigenvalues of matrix "A". To do it, solve the characteristic equation.


"\det(A-\lambda I) = 0"


"\det(A-\lambda I) = \begin{vmatrix}
 -1-\lambda & \frac{1}{2} & \frac{1}{2} \\
 \frac{1}{2} & 1-\lambda & -3 \\
 \frac{1}{2} & -3 & 1-\lambda
\end{vmatrix}"


"= (-1-\lambda)\begin{vmatrix}
 1 - \lambda & -3 \\
 -3 & 1- \lambda
\end{vmatrix} - \frac{1}{2} \begin{vmatrix}
 \frac{1}{2} & -3 \\
 \frac{1}{2} & 1- \lambda
\end{vmatrix} + \frac{1}{2} \begin{vmatrix}
 \frac{1}{2} & 1 - \lambda \\
 \frac{1}{2} & -3
\end{vmatrix}"


"= (-1- \lambda)[(1- \lambda)^2 - 9] - \frac{1}{2}[\frac{1}{2}(1- \lambda) + \frac{3}{2}] + \frac{1}{2}[-\frac{3}{2} - \frac{1}{2}(1- \lambda)]"


"= (-1- \lambda)(-2- \lambda)(4- \lambda) - \frac{1}{4}(4- \lambda) - \frac{1}{4}(4 - \lambda)"


"= (4 - \lambda)( \lambda^2 + 3 \lambda +\frac{3}{2})"


Using the formula for discriminant for the second term, obtain


"(4 - \lambda)(\lambda + \frac{3}{2} + \frac{\sqrt{3}}{2})(\lambda + \frac{3}{2} - \frac{\sqrt{3}}{2}) = 0"


Therefore, the eigenvalues, in the decreasing order, are


"\lambda_{1} = 4, \lambda_{2} = -\frac{3}{2} + \frac{\sqrt{3}}{2}, \lambda_{3} = -\frac{3}{2} - \frac{\sqrt{3}}{2}"


Thus, the orthogonal canonical reduction of "Q" is


"Q = 4x_{1}^{2} + ( -\frac{3}{2} + \frac{\sqrt{3}}{2})y_{1}^{2} + ( -\frac{3}{2} - \frac{\sqrt{3}}{2})z_{1}^{2} \\ \\
\quad=4x_{1}^{2} - (\frac{3}{2} - \frac{\sqrt{3}}{2})y_{1}^{2} - ( \frac{3}{2} + \frac{\sqrt{3}}{2})z_{1}^{2}"


where "x_{1}, y_{1}, z_{1}" are the new coordinates.


If "\lambda = 4", then


"\begin{bmatrix}
 -1-4 & \frac{1}{2} & \frac{1}{2} \\
 \frac{1}{2} & 1-4 & -3 \\
 \frac{1}{2} & -3 & 1-4
\end{bmatrix}\begin{bmatrix}
 x \\
 y \\
z
\end{bmatrix} =\begin{bmatrix}
 0 \\
 0 \\
0
\end{bmatrix}"


"\begin{bmatrix}
 -5 & \frac{1}{2} & \frac{1}{2} \\
 \frac{1}{2} & -3 & -3 \\
 \frac{1}{2} & -3 & -3
\end{bmatrix}\begin{bmatrix}
 x \\
 y \\
z
\end{bmatrix} =\begin{bmatrix}
 0 \\
 0 \\
0
\end{bmatrix}"


"\begin{bmatrix}
 x \\
 y \\
z
\end{bmatrix} =\begin{bmatrix}
 0 \\
 -1 \\
1
\end{bmatrix}"


The first principal axis is


"p_{1} = \frac{1}{\sqrt{0^2 + (-1)^{2}+1^2}} \begin{bmatrix}
 0 \\
 -1 \\
1
\end{bmatrix} = \frac{1}{\sqrt{2}} \begin{bmatrix}
 0 \\
 -1 \\
1
\end{bmatrix}"


If "\lambda = -\frac{3}{2} + \frac{\sqrt{3}}{2}", then


"\begin{bmatrix}
 -1 +\frac{3}{2} - \frac{\sqrt{3}}{2} & \frac{1}{2} & \frac{1}{2} \\
 \frac{1}{2} & 1+\frac{3}{2} - \frac{\sqrt{3}}{2} & -3 \\
 \frac{1}{2} & -3 & 1+\frac{3}{2} - \frac{\sqrt{3}}{2} 
\end{bmatrix}\begin{bmatrix}
 x \\
 y \\
z
\end{bmatrix} =\begin{bmatrix}
 0 \\
 0 \\
0
\end{bmatrix}"



"\begin{bmatrix}
 \frac{1}{2} - \frac{\sqrt{3}}{2} & \frac{1}{2} & \frac{1}{2} \\
 \frac{1}{2} & \frac{5}{2} - \frac{\sqrt{3}}{2} & -3 \\
 \frac{1}{2} & -3 & \frac{5}{2} - \frac{\sqrt{3}}{2}
\end{bmatrix}\begin{bmatrix}
 x \\
 y \\
z
\end{bmatrix} =\begin{bmatrix}
 0 \\
 0 \\
0
\end{bmatrix}"


"\begin{bmatrix}
 x \\
 y \\
z
\end{bmatrix} =\begin{bmatrix}
 1+\sqrt{3} \\
 1 \\
1
\end{bmatrix}"


The second principal axis is


"p_{2} = \frac{1}{\sqrt{(1+\sqrt{3})^2 + 1^{2}+1^2}} \begin{bmatrix}
 1+\sqrt{3} \\
 1 \\
1
\end{bmatrix} \\
\quad = \frac{1}{\sqrt{6+2\sqrt{3}}} \begin{bmatrix}
 1+\sqrt{3} \\
 1 \\
1
\end{bmatrix}"


If "\lambda = -\frac{3}{2} - \frac{\sqrt{3}}{2}", then


"\begin{bmatrix}
 -1 +\frac{3}{2} + \frac{\sqrt{3}}{2} & \frac{1}{2} & \frac{1}{2} \\
 \frac{1}{2} & 1+\frac{3}{2} + \frac{\sqrt{3}}{2} & -3 \\
 \frac{1}{2} & -3 & 1+\frac{3}{2} + \frac{\sqrt{3}}{2} 
\end{bmatrix}\begin{bmatrix}
 x \\
 y \\
z
\end{bmatrix} =\begin{bmatrix}
 0 \\
 0 \\
0
\end{bmatrix}"



"\begin{bmatrix}
 \frac{1}{2} -+\frac{\sqrt{3}}{2} & \frac{1}{2} & \frac{1}{2} \\
 \frac{1}{2} & \frac{5}{2} + \frac{\sqrt{3}}{2} & -3 \\
 \frac{1}{2} & -3 & \frac{5}{2} + \frac{\sqrt{3}}{2}
\end{bmatrix}\begin{bmatrix}
 x \\
 y \\
z
\end{bmatrix} =\begin{bmatrix}
 0 \\
 0 \\
0
\end{bmatrix}"



"\begin{bmatrix}
 x \\
 y \\
z
\end{bmatrix} =\begin{bmatrix}
 1-\sqrt{3} \\
 1 \\
1
\end{bmatrix}"


The third principal axis is


"p_{3} = \frac{1}{\sqrt{(1-\sqrt{3})^2 + 1^{2}+1^2}} \begin{bmatrix}
 1-\sqrt{3} \\
 1 \\
1
\end{bmatrix} \\

\quad= \frac{1}{\sqrt{6-2\sqrt{3}}} \begin{bmatrix}
 1-\sqrt{3} \\
 1 \\
1
\end{bmatrix}"


 The change of basis needed to convert "Q" into the orthogonal canonical reduction is given by


"\begin{bmatrix}
 x \\
 y \\
z
\end{bmatrix} =\begin{bmatrix}
 0 & \frac{1 + \sqrt{3}}{\sqrt{6+2\sqrt{3}}} & \frac{1 - \sqrt{3}}{\sqrt{6+-\sqrt{3}}} \\
 -\frac{1}{\sqrt{2}} & \frac{1}{\sqrt{6+2\sqrt{3}}} & \frac{1}{\sqrt{6-2\sqrt{3}}} \\
 \frac{1}{\sqrt{2}} & \frac{1}{\sqrt{6+2\sqrt{3}}} & \frac{1}{\sqrt{6-2\sqrt{3}}}
\end{bmatrix} \begin{bmatrix}
 x_1 \\
 y_1 \\
z_1
\end{bmatrix}"